erin drove home at an average speed of 45 mph. she returned home in the evening along the same route and average 30 mph. if erin spent a total of 1 hour commuting to and from work, how many miles did erin drive to work in the morning?
Found 2 solutions by stanbon, RAY100:
Answer by stanbon(48558)

erin drove home at an average speed of 45 mph.
DATA:
rate = 45 mph ; distance = x miles ; time = x/45 hrs
---------------------------------------------------------
she returned home in the evening along the same route and average 30 mph.
DATA:
rate = 30 mph ; distance = x miles ; time = x/30 hrs
-------------------------------------------------------
if erin spent a total of 1 hour commuting to and from work, how many miles did erin drive to work in the morning?
-----------
Equation:
time to + time from = 1 hr
x/45 + x/30 = 1
(30x + 45x) = 30*45
75x = 1350
x = 18 miles (distance to work in the morning)
======================================================
Cheers,
Stan H.

Answer by RAY100(1637)

Let vel=Distance/Time
let d = distance to work
t1=d/vel=d/45
t2= d/30
t1+t2=1,,,,given
d/45+d/30=1
let LCD =90
(2d +3d)/90 = 1
5d=90
d=18
check
18/45 +18/30 =1
.4 + .6=1,,,,ok